One-tailed test, o? is known, a = 0.05, n= 34


1
Expert's answer
2022-05-03T11:36:48-0400

"n = 34, α = 0.05, \sigma" is known.

One-tailed test

Given "df=n-1=33" degrees of freedom. The critical "t-" value for one-tailed, "df=33" degrees of freedom, and "\alpha=0.05" is


"t_c=1.69236, right-tailed"

"t_c=-1.69236, left-tailed"